# Discriminative Adversarial Unlearning

### 10 Feb 2024 v1

@@ Abstract
We introduce a novel machine unlearning framework founded upon the established principles of the min-max optimization paradigm. We capitalize on the capabilities of strong Membership Inference Attacks (MIA) to facilitate the unlearning of specific samples from a trained model. We consider the scenario of two networks, the attacker $\mathbf{A}$ and the trained defender $\mathbf{D}$ pitted against each other in an adversarial objective, wherein the attacker aims at teasing out the information of the data to be unlearned in order to infer membership, and the defender unlearns to defend the network against the attack, whilst preserving its general performance. The algorithm can be trained end-to-end using backpropagation, following the well known iterative min-max approach in updating the attacker and the defender. We additionally incorporate a self-supervised objective effectively addressing the feature space discrepancies between the forget set and the validation set, enhancing unlearning performance. Our proposed algorithm closely approximates the ideal benchmark of retraining from scratch for both random sample forgetting and class-wise forgetting schemes on standard machine-unlearning datasets. Specifically, on the class unlearning scheme, the method demonstrates near-optimal performance and comprehensively overcomes known methods over the random sample forgetting scheme across all metrics and multiple network pruning strategies.
@@ 1 Introduction
With over \$4 billion paid in settlements over privacy concerns by Big Tech firms since the enforcement of the General Data Protection Regulation [16] in 2018  [^1] Source: Data Privacy Manager, and more than 70% of citizens from nine major countries expressing support for increased governmental intervention in Big Tech [^2] Source: Amnesty International, significant concerns arise regarding individual privacy and the potential erosion of trust between human users and AI systems. Owing to their widespread incorporation, digital systems now function as vast repositories of personal data, creating a comprehensive digital footprint that reflects a broad spectrum of online behaviors, interactions, and communication patterns [45]. This footprint is evident across various forms of user-generated content, including product reviews, blog posts, social media activities, and contributions to collaborative platforms such as Wikipedia [46, 8]. The application of artificial intelligence (AI) in areas such as targeted advertising and personalized product placement, along with the emergence of potentially harmful technologies like deepfakes, further underscores a critical concern regarding the lack of transparency and control over personal data, highlighting the need for protective measures in the management and utilization of personal information [60, 1]. The recent legal developments echo these concerns with a notable development being, ”the right to be forgotten,” [14, 53] which requires the deletion of personal data from both digital databases and machine learning models at an individual’s request. However, retraining AI models to accommodate each deletion request is impractical, given the potential cost of millions of dollars, coupled with substantial environmental impact.
@@
In response to these challenges, there has been a notable surge in the development of algorithmic unlearning strategies [57, 63, 30, 10, 37, 38, 9]. These strategies aim to modify existing models, allowing them to selectively ’unlearn’ specific samples from their training data. The approaches offer viable alternatives to complete retraining from scratch, providing cost-effective solutions that enhance data security and privacy while preserving the efficacy of the models. Despite their significant effectiveness, a majority of these measures [38, 42, 18, 28, 29] primarily address the problem in the parametric space. They introduce modifications within the pre-training objectives [37] or the pre-processing phase [40, 20] of trained models, often entailing performance-cost trade-offs due to the evaluation of hessian approximations or proposals for complete overhauls of training procedures.
@@
Membership Inference Attacks (MIA), on the other hand as a field of study, has concurrently bloomed [50, 48, 44, 36, 51, 61] exposing significant privacy related threats to trained models. For a trained model, MIAs are tasked with identification of inference of patterns in the model’s behaviour due to samples arising from its training set and other sets. Using these patterns, a successful attacker may avail themselves with personally identifying information [56, 17, 55, 19] posing significant threats to the use of AI towards general welfare.
@@
  • We innovate by repurposing the Barlow Twins objective [62] for unlearning, yielding a substantial performance boost. This repurposing opens avenues for applying self-supervised learning research to provide practical solutions in machine unlearning.
@@ 2.1 Machine Unlearning
Machine unlearning endeavors to mitigate the influence of specific data points or classes from a trained ML model, primarily addressing privacy concerns in accordance with regulations [47, 31]. Recent developments in the field has led to a taxonomy for methods broadly classifying the objectives as follows.
@@
Retraining the model from the scratch on the data filtering out forgetting dataset is the most straightforward and effective manner to update ML models for unlearning, ensuring complete deletion of undesirable samples or aspects from a pre-trained model. Cao and Yang [6] explore this mode of unlearning over Naive Bayes classifiers whereas Ginart et al. [22] implement deletion algorithms for $k\text{-}means$ clustering, rendering them unscalable to deep neural networks. Additionally, recent efforts [5] also encourage training on shards of data, subsequently incorporating aggregation, effectively alleviating the expensive nature of training a large model from scratch. However, the measures signify impracticality in the face of large queries for deletion, requiring inordinate amounts of compute especially for models wherein the training cost estimates exceed millions of dollars.
@@
Acknowledgement of the said infeasibility of training large models from scratch has therefore led to an upsurge in palliative measures to selectively eradicate information from trained models. Model finetuning as an approach [58, 23] aims to finetune the model on the retain set of data to induce catastrophic forgetting within trained models. Graves et al. [25], Thudi et al. [54] on the other hand propose retraining the model on corrupted labels for sensitive data providing for a relatively cost-effective measure to induce unlearning. Hessian matrix approximation largely characterizes some of the works [42]. Use of fisher information matrices [4, 39] allows for a direct albeit expensive estimation of the expectation of the hessian due to $\log p(y|x,\theta)$, allowing for strategies pertinent to noising/masking if the said estimation for the data to be forgotten. Fisher forgetting [4, 39] broadly aims at a similar estimation of the hessians using FIMs in order to conduct informed corruption of samples in the image space. Model decisions on the other hand may lend to some avenues for manipulation [10, 57] in order to induce forgetting using statistical distance based metrics such as $KL$ divergence. Unlearning measures have also experienced a surge in interest within the generative models space and in Large Language Models (LLMs). In the generative model domain [28, 29, 18], efforts are directed towards inhibiting diffusion or Generative Adversarial Network (GAN)-based models from generating questionable content. Model editing methods [13, 21, 43] aim to edit factual information encoded within pre-trained LLMs. However, these approaches predominantly focus on selective influential aspects of the trained model in isolation, overlooking the potential for incorporating separate networks to expedite the unlearning process in the model under consideration. Furthermore, in order to incorporate expensive measures for estimation of second order information, significant approximations are formulated which determine the efficiency and performances of the methods.
@@ 2.2 Membership Inference Attacks (MIA)
The purpose of a MIA is to uncover from a trained model (target), information revealing whether a sample was involved in the training process of the target model. Assuming black box conditions, the attack model trains itself to make binary predictions using the outputs of the target, inferring belonging or lack thereof of samples in the training set of target. This formulation was first proposed by Shokri et al. [50], wherein they create multiple shadow models to replicate the behavior of a target model. The shadow models are then utilized to generate data for training a binary classifier to determine membership. Building upon this concept, Salem et al. [48] enhanced the efficacy of the attack by employing a single shadow model. Refinements to the method were achieved through the integration of supplementary discriminative information [44], which included considerations of labels. Additional contributions [36, 59, 52, 51] introduced evaluative metrics where the membership status of a record is directly ascertained by comparing predefined thresholds based on metrics such as prediction confidences, entropy, or modified entropy of the record. Song and Mittal demonstrated that, by establishing a class-dependent threshold, the metric-based classifier could achieve comparable or even superior inference performance compared to the neural network-based classifier [51]. In this work we adopt the recent work of Yuan and Zhang [61] towards proposing an effective mechanism for MIA on pruned models. Briefly, the work proposes a transformer based architecture that incorporates multiple sources of information about the target model’s inferential patterns in order to effectively conduct MIA against the model.
@@ 3.1 The Attacker
Consider a target defender model $\mathbf{D}$, which is trained on the training set $\mathcal{D}$. The objective of the attacker model $\mathbf{A}$, is to deduce a binary prediction indicating the presence or absence of a specific sample in the training set. Typically, the scenario is characterized as a black-box setting, where the attacker’s access is constrained to the outputs of the defender model. The attacker model is trained by utilizing the defender’s inferences on samples originating from both the set of forgotten samples $\mathcal{D}_{f}$ and the validation set $\mathcal{D}_{v}$, with the distinction that the former was used in the training process of the defender, while the latter was not. As we shall see later, it is strategically advantageous to employ a potent attacker that assimilates substantial information to achieve accurate inferences regarding membership. Accordingly, we consider the recent formulation proposed by Yuan and Zhang [61]. Briefly, this attacker is designed to process three crucial pieces of information for each sample:
@@ 3.3 Adversarial Unlearning
where $\alpha$ modulates the influence of the attacker’s objective on the defender network which also optimizes for a $K$ class classification problem. To facilitate the training of the networks with the specified objective, we employ the iterative updating scheme common to the min-max optimization frameworks for deep network learning [24, 2, 26], for both the defender and the attacker. The trained defender is provided with samples from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$ on which it conducts inference. The attacker subsequently uses these inferences to generate its inputs, as outlined in Section 3.1. During this process, the attacker acquires knowledge of the inference patterns exhibited by the defender and endeavors to distinguish between the defender’s outputs for samples from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$, by maximizing a binary classification objective. This discrimination task is relatively easier for the attacker in the early stages of the algorithm. As the training progresses, the defender becomes aware of the attacker’s motives and consequently adjusts its own learning objectives to defend against the attack, minimizing the attacker’s objective, leading to a decline in its performance on the set of retained samples $\mathcal{D}_{r}$. In response to this degradation, the defender optimizes for learning to classify on $\mathcal{D}_{r}$ to preserve its general capacity for classification on $\mathcal{D}$. This adaptive learning process reflects the dynamic interplay between the defender and the attacker during training, ultimately influencing the defender’s strategies to balance classification performance on the retained set and resistance against the MIA.
One may however consider making practical adjustments to the objective, in consideration of the problem of weaker gradients as identified in the work of Generative Adversatial Nets (GAN) Goodfellow et al. [24] and instead maximize $\mathbb{E}_{x_{f}\sim\mathcal{D}_{f},x_{v}\sim\mathcal{D}_{v}}[\log(\mathbf{A}(1-\mathbf{D}(x_{f}))+\log(\mathbf{A}(\mathbf{D}(x_{v})))]$ to empirically adjust for the differential capacities of the attacker and the defender networks to counter each other’s efforts. Additionally, one may also incorporate improvements over the GAN algorithm [2, 26, 34] in order to ensure better convergence. Incorporation of objectives pertinent to unlearning such as gradient ascent using a negative of the loss function for $\mathcal{D}_{f}$ [25, 54], or fisher forgetting through additive gaussian noise [4, 32] is also feasible through relatively nominal adjustments to the algorithm.
@@ 3.4 Self-supervised Regularization
The min-max objective strives to render the outputs of $\mathbf{D}$ indistinguishable to a proficient attacker within the output space. Nevertheless, the defender network tends to overcompensate by emphasizing exclusively on the final output layer, aiming to enhance its competence against the attacker, and the feature space discrepancies between $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$ remain overlooked. In order to address these discrepancies we introduce a self supervised regularization term that enables the defender to inherit the task of forgetting samples within the network’s parameters and eases the burden on the final layer of the network. We motivate this objective through the notion that the features generated for the forget set $\mathcal{D}_{f}$ must on average be similar to those arising from the validation set $\mathcal{D}_{v}$ and therefore the test set $\mathcal{D}_{t}$. In our application we consider the feature space redundancy reduction principle utilized by Zbontar et al. [62] that aims to enforce invariance between features of samples that in principle must be similar. The defender $\mathbf{D}$ essentially consisting of the encoder and the FC layer $\mathbf{D}(.)=\mathbf{D}_{fc}(\mathbf{D}_{enc}(.))$, upon processing the input samples, outputs mean-centered features $\{z_{b}:z_{b}=\mathbf{D}_{enc}(x_{b}),z_{b}\in\mathbb{R}^{D}\}$, where $b$ indexes the batch. Subsequently, for a cross-correlation matrix $\mathcal{C}$ computed between features from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$, defined as,
@@
where $\lambda$ was set to $-35e\text{-}$ 3 as recommended in Zbontar et al. [62]. The regularization plays a crucial role in augmenting the overall performance of the algorithm, as elucidated through ablation studies in Appendix Section 4.5. However, one may surmise several other variants of the self-supervised objective [3, 7]. Additionally, contrastive objectives [11, 49, 15] that enable similarity between $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$ and distance between $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, may also be incorporated as regularizers. One may also leverage class label information to motivate self-supervised objectives within the classes that the samples belong to. Furthermore, one may incorporate a separate projector network over the defender in order to implement this regularization as is commonly practiced in SSL literature. The overall objective that we utilize in order to train the framework through gradient descent is formulated as:
@@
Table 1: Cifar-10, Cifar-100 comparison. We evaluate the unlearning performance on Cifar datasets. The top two tables encompass the unlearning performances on Cifar-10 for the methods listed for random and class-wise schemes respectively. Evaluation metrics from section 4 are employed, including the averaged disparity of performances relative to retraining on the retain set $\mathcal{D}r$. Five separate seeds are utilized for each method, and the reported mean and standard deviation are presented in the format $a{\pm b}$. Additionally, unlearning is performed on $95\%$ sparse models for each method, employing OMP [41] to enforce sparsity. The relative performances are compared against the ”retrain” standard, and the average disparity provides an overarching assessment of all algorithms across the employed metrics. (UA: Unlearning Acc., RA: Retaining Acc., TA: Test Acc.)

### 13 Feb 2024 v2

@@ Abstract
We introduce a novel machine unlearning framework founded upon the established principles of the min-max optimization paradigm. We capitalize on the capabilities of strong Membership Inference Attacks (MIA) to facilitate the unlearning of specific samples from a trained model. We consider the scenario of two networks, the attacker $\mathbf{A}$ and the trained defender $\mathbf{D}$ pitted against each other in an adversarial objective, wherein the attacker aims at teasing out the information of the data to be unlearned in order to infer membership, and the defender unlearns to defend the network against the attack, whilst preserving its general performance. The algorithm can be trained end-to-end using backpropagation, following the well known iterative min-max approach in updating the attacker and the defender. We additionally incorporate a self-supervised objective effectively addressing the feature space discrepancies between the forget set and the validation set, enhancing unlearning performance. Our proposed algorithm closely approximates the ideal benchmark of retraining from scratch for both random sample forgetting and class-wise forgetting schemes on standard machine-unlearning datasets. Specifically, on the class unlearning scheme, the method demonstrates near-optimal performance and comprehensively overcomes known methods over the random sample forgetting scheme across all metrics and multiple network pruning strategies. Code is available at: Link.
Machine Learning, ICML
@@ 1 Introduction
With over \$4 billion paid in settlements over privacy concerns by Big Tech firms since the enforcement of the General Data Protection Regulation (European Parliament and Council of the European Union, 2016) in 2018  [^1] Source: Data Privacy Manager, and more than 70% of citizens from nine major countries expressing support for increased governmental intervention in Big Tech [^2] Source: Amnesty International, significant concerns arise regarding individual privacy and the potential erosion of trust between human users and AI systems. Owing to their widespread incorporation, digital systems now function as vast repositories of personal data, creating a comprehensive digital footprint that reflects a broad spectrum of online behaviors, interactions, and communication patterns (Nguyên, 2019). This footprint is evident across various forms of user-generated content, including product reviews, blog posts, social media activities, and contributions to collaborative platforms such as Wikipedia (Nguyen et al., 2021; Carpentier et al., 2021). The application of artificial intelligence (AI) in areas such as targeted advertising and personalized product placement, along with the emergence of potentially harmful technologies like deepfakes, further underscores a critical concern regarding the lack of transparency and control over personal data, highlighting the need for protective measures in the management and utilization of personal information (Yevseiev et al., 2021; Al-Harrasi et al., 2023). The recent legal developments echo these concerns with a notable development being, ”the right to be forgotten,” (Dang, 2021; Tankard, 2016) which requires the deletion of personal data from both digital databases and machine learning models at an individual’s request. However, retraining AI models to accommodate each deletion request is impractical, given the potential cost of millions of dollars, coupled with substantial environmental impact.
@@
In response to these challenges, there has been a notable surge in the development of algorithmic unlearning strategies (Wang et al., 2023; Zhang et al., 2023; Hoang et al., 2024; Chen et al., 2023; Lin et al., 2023; Liu et al., 2023a; Che et al., 2023). These strategies aim to modify existing models, allowing them to selectively ’unlearn’ specific samples from their training data. The approaches offer viable alternatives to complete retraining from scratch, providing cost-effective solutions that enhance data security and privacy while preserving the efficacy of the models. Despite their significant effectiveness, a majority of these measures (Liu et al., 2023a; Mehta et al., 2022; Gandikota et al., 2023; Heng and Soh, 2023a, b) primarily address the problem in the parametric space. They introduce modifications within the pre-training objectives (Lin et al., 2023) or the pre-processing phase (Liu et al., 2017; García-Pablos et al., 2020) of trained models, often entailing performance-cost trade-offs due to the evaluation of hessian approximations or proposals for complete overhauls of training procedures.
@@
Membership Inference Attacks (MIA), on the other hand as a field of study, has concurrently bloomed (Shokri et al., 2017; Salem et al., 2018; Nasr et al., 2018; Leino and Fredrikson, 2020; Song and Mittal, 2021; Yuan and Zhang, 2022) exposing significant privacy related threats to trained models. For a trained model, MIAs are tasked with identification of inference of patterns in the model’s behaviour due to samples arising from its training set and other sets. Using these patterns, a successful attacker may avail themselves with personally identifying information (Veale et al., 2018; Fredrikson et al., 2015; Tramèr et al., 2016; Ganju et al., 2018) posing significant threats to the use of AI towards general welfare.
@@
  • We innovate by repurposing the Barlow Twins objective (Zbontar et al., 2021) for unlearning, yielding a substantial performance boost. This repurposing opens avenues for applying self-supervised learning research to provide practical solutions in machine unlearning.
@@ 2.1 Machine Unlearning
Machine unlearning endeavors to mitigate the influence of specific data points or classes from a trained ML model, primarily addressing privacy concerns in accordance with regulations (Rosen, 2011; Hoofnagle et al., 2019). Recent developments in the field has led to a taxonomy for methods broadly classifying the objectives as follows.
@@
Retraining the model from the scratch on the data filtering out forgetting dataset is the most straightforward and effective manner to update ML models for unlearning, ensuring complete deletion of undesirable samples or aspects from a pre-trained model. Cao and Yang (2015) explore this mode of unlearning over Naive Bayes classifiers whereas Ginart et al. (2019) implement deletion algorithms for $k\text{-}means$ clustering, rendering them unscalable to deep neural networks. Additionally, recent efforts (Bourtoule et al., 2021) also encourage training on shards of data, subsequently incorporating aggregation, effectively alleviating the expensive nature of training a large model from scratch. However, the measures signify impracticality in the face of large queries for deletion, requiring inordinate amounts of compute especially for models wherein the training cost estimates exceed millions of dollars.
@@
Acknowledgement of the said infeasibility of training large models from scratch has therefore led to an upsurge in palliative measures to selectively eradicate information from trained models. Model finetuning as an approach (Warnecke et al., 2021; Golatkar et al., 2020) aims to finetune the model on the retain set of data to induce catastrophic forgetting within trained models. Graves et al. (2021); Thudi et al. (2022) on the other hand propose retraining the model on corrupted labels for sensitive data providing for a relatively cost-effective measure to induce unlearning. Hessian matrix approximation largely characterizes some of the works (Mehta et al., 2022). Use of fisher information matrices (Becker and Liebig, 2022; Liu et al., 2023b) allows for a direct albeit expensive estimation of the expectation of the hessian due to $\log p(y|x,\theta)$, allowing for strategies pertinent to noising/masking if the said estimation for the data to be forgotten. Fisher forgetting (Becker and Liebig, 2022; Liu et al., 2023b) broadly aims at a similar estimation of the hessians using FIMs in order to conduct informed corruption of samples in the image space. Model decisions on the other hand may lend to some avenues for manipulation (Chen et al., 2023; Wang et al., 2023) in order to induce forgetting using statistical distance based metrics such as $KL$ divergence. Unlearning measures have also experienced a surge in interest within the generative models space and in Large Language Models (LLMs). In the generative model domain (Heng and Soh, 2023a, b; Gandikota et al., 2023), efforts are directed towards inhibiting diffusion or Generative Adversarial Network (GAN)-based models from generating questionable content. Model editing methods (Dai et al., 2021; Geva et al., 2020; Meng et al., 2022) aim to edit factual information encoded within pre-trained LLMs. However, these approaches predominantly focus on selective influential aspects of the trained model in isolation, overlooking the potential for incorporating separate networks to expedite the unlearning process in the model under consideration. Furthermore, in order to incorporate expensive measures for estimation of second order information, significant approximations are formulated which determine the efficiency and performances of the methods.
@@ 2.2 Membership Inference Attacks (MIA)
The purpose of a MIA is to uncover from a trained model (target), information revealing whether a sample was involved in the training process of the target model. Assuming black box conditions, the attack model trains itself to make binary predictions using the outputs of the target, inferring belonging or lack thereof of samples in the training set of target. This formulation was first proposed by Shokri et al. (2017), wherein they create multiple shadow models to replicate the behavior of a target model. The shadow models are then utilized to generate data for training a binary classifier to determine membership. Building upon this concept, Salem et al. (2018) enhanced the efficacy of the attack by employing a single shadow model. Refinements to the method were achieved through the integration of supplementary discriminative information (Nasr et al., 2018), which included considerations of labels. Additional contributions (Leino and Fredrikson, 2020; Yeom et al., 2018; Song et al., 2019; Song and Mittal, 2021) introduced evaluative metrics where the membership status of a record is directly ascertained by comparing predefined thresholds based on metrics such as prediction confidences, entropy, or modified entropy of the record. Song and Mittal demonstrated that, by establishing a class-dependent threshold, the metric-based classifier could achieve comparable or even superior inference performance compared to the neural network-based classifier (Song and Mittal, 2021). In this work we adopt the recent work of Yuan and Zhang (2022) towards proposing an effective mechanism for MIA on pruned models. Briefly, the work proposes a transformer based architecture that incorporates multiple sources of information about the target model’s inferential patterns in order to effectively conduct MIA against the model.
@@ 3.1 The Attacker
Consider a target defender model $\mathbf{D}$, which is trained on the training set $\mathcal{D}$. The objective of the attacker model $\mathbf{A}$, is to deduce a binary prediction indicating the presence or absence of a specific sample in the training set. Typically, the scenario is characterized as a black-box setting, where the attacker’s access is constrained to the outputs of the defender model. The attacker model is trained by utilizing the defender’s inferences on samples originating from both the set of forgotten samples $\mathcal{D}_{f}$ and the validation set $\mathcal{D}_{v}$, with the distinction that the former was used in the training process of the defender, while the latter was not. As we shall see later, it is strategically advantageous to employ a potent attacker that assimilates substantial information to achieve accurate inferences regarding membership. Accordingly, we consider the recent formulation proposed by Yuan and Zhang (2022). Briefly, this attacker is designed to process three crucial pieces of information for each sample:
@@ 3.3 Adversarial Unlearning
where $\alpha$ modulates the influence of the attacker’s objective on the defender network which also optimizes for a $K$ class classification problem. To facilitate the training of the networks with the specified objective, we employ the iterative updating scheme common to the min-max optimization frameworks for deep network learning (Goodfellow et al., 2020; Arjovsky et al., 2017; Gulrajani et al., 2017), for both the defender and the attacker. The trained defender is provided with samples from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$ on which it conducts inference. The attacker subsequently uses these inferences to generate its inputs, as outlined in Section 3.1. During this process, the attacker acquires knowledge of the inference patterns exhibited by the defender and endeavors to distinguish between the defender’s outputs for samples from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$, by maximizing a binary classification objective. This discrimination task is relatively easier for the attacker in the early stages of the algorithm. As the training progresses, the defender becomes aware of the attacker’s motives and consequently adjusts its own learning objectives to defend against the attack, minimizing the attacker’s objective, leading to a decline in its performance on the set of retained samples $\mathcal{D}_{r}$. In response to this degradation, the defender optimizes for learning to classify on $\mathcal{D}_{r}$ to preserve its general capacity for classification on $\mathcal{D}$. This adaptive learning process reflects the dynamic interplay between the defender and the attacker during training, ultimately influencing the defender’s strategies to balance classification performance on the retained set and resistance against the MIA.
One may however consider making practical adjustments to the objective, in consideration of the problem of weaker gradients as identified in the work of Generative Adversatial Nets (GAN) Goodfellow et al. (2020) and instead maximize $\mathbb{E}_{x_{f}\sim\mathcal{D}_{f},x_{v}\sim\mathcal{D}_{v}}[\log(\mathbf{A}(1-\mathbf{D}(x_{f}))+\log(\mathbf{A}(\mathbf{D}(x_{v})))]$ to empirically adjust for the differential capacities of the attacker and the defender networks to counter each other’s efforts. Additionally, one may also incorporate improvements over the GAN algorithm (Arjovsky et al., 2017; Gulrajani et al., 2017; Kodali et al., 2017) in order to ensure better convergence. Incorporation of objectives pertinent to unlearning such as gradient ascent using a negative of the loss function for $\mathcal{D}_{f}$ (Graves et al., 2021; Thudi et al., 2022), or fisher forgetting through additive gaussian noise (Becker and Liebig, 2022; Izzo et al., 2021) is also feasible through relatively nominal adjustments to the algorithm.
@@ 3.4 Self-supervised Regularization
The min-max objective strives to render the outputs of $\mathbf{D}$ indistinguishable to a proficient attacker within the output space. Nevertheless, the defender network tends to overcompensate by emphasizing exclusively on the final output layer, aiming to enhance its competence against the attacker, and the feature space discrepancies between $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$ remain overlooked. In order to address these discrepancies we introduce a self supervised regularization term that enables the defender to inherit the task of forgetting samples within the network’s parameters and eases the burden on the final layer of the network. We motivate this objective through the notion that the features generated for the forget set $\mathcal{D}_{f}$ must on average be similar to those arising from the validation set $\mathcal{D}_{v}$ and therefore the test set $\mathcal{D}_{t}$. In our application we consider the feature space redundancy reduction principle utilized by Zbontar et al. (2021) that aims to enforce invariance between features of samples that in principle must be similar. The defender $\mathbf{D}$ essentially consisting of the encoder and the FC layer $\mathbf{D}(.)=\mathbf{D}_{fc}(\mathbf{D}_{enc}(.))$, upon processing the input samples, outputs mean-centered features $\{z_{b}:z_{b}=\mathbf{D}_{enc}(x_{b}),z_{b}\in\mathbb{R}^{D}\}$, where $b$ indexes the batch. Subsequently, for a cross-correlation matrix $\mathcal{C}$ computed between features from $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$, defined as,
@@
where $\lambda$ was set to $-35e\text{-}$ 3 as recommended in Zbontar et al. (2021). The regularization plays a crucial role in augmenting the overall performance of the algorithm, as elucidated through ablation studies in Appendix Section 4.5. However, one may surmise several other variants of the self-supervised objective (Bardes et al., 2021; Caron et al., 2021). Additionally, contrastive objectives (Chen et al., 2020; Sharma et al., 2024; Dwibedi et al., 2021) that enable similarity between $\mathcal{D}_{f}$ and $\mathcal{D}_{v}$ and distance between $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, may also be incorporated as regularizers. One may also leverage class label information to motivate self-supervised objectives within the classes that the samples belong to. Furthermore, one may incorporate a separate projector network over the defender in order to implement this regularization as is commonly practiced in SSL literature. The overall objective that we utilize in order to train the framework through gradient descent is formulated as:
@@
Table 1: Cifar-10, Cifar-100 comparison. We evaluate the unlearning performance on Cifar datasets. The top two tables encompass the unlearning performances on Cifar-10 for the methods listed for random and class-wise schemes respectively. Evaluation metrics from section 4 are employed, including the averaged disparity of performances relative to retraining on the retain set $\mathcal{D}r$. Five separate seeds are utilized for each method, and the reported mean and standard deviation are presented in the format $a{\pm b}$. Additionally, unlearning is performed on $95\%$ sparse models for each method, employing OMP (Ma et al., 2021) to enforce sparsity. The relative performances are compared against the ”retrain” standard, and the average disparity provides an overarching assessment of all algorithms across the employed metrics. (UA: Unlearning Acc., RA: Retaining Acc., TA: Test Acc.)